OpenStudy (anonymous):

Let f(x) = x2 + 3x – 4 and g(x) = x + 5. Find f(x) • g(x). x3 + 3x2 + 16x – 20 x3 + 5x2 + 14x – 20 x3 + 8x2 + 11x – 20 x3 + 9x2 + 19x – 20

OpenStudy (anonymous):

c

OpenStudy (anonymous):

f(x)g(x)=(x2+3x-4)(x+5)=x3+5x2+3x2+15x-4x-20=x3+8x2+11x-20, Thus c

OpenStudy (anonymous):